### v3

@@ 5 Conclusion
This paper demonstrates the feasibility of extending the optimization of Stefan problems to scenarios where the flow in the fluid phase is significant and cannot be ignored. By treating the velocity field as a checkpoint variable, we were able to derive an incomplete continuous adjoint formulation for the two-phase problem. Our results indicate that the adjoint-based gradient descent method effectively controls the shape of melting fronts in this highly nonlinear system. Particularly, the number of function evaluations is greatly reduced with respect to derivative-free algorithms. This approach shows great potential for applications in more complex and dynamic melting processes. Looking forward, future work could explore more advanced actuator models, especially in non-periodic domains, to further enhance the optimization framework and broaden its applicability.
@@
T.F. and T.S. conceived the study. T.F. performed the numerical simulations with feedback from all authors. T.F. wrote the original draft with feedback from all authors. T.S. supervised the study.

### v4

@@ 5 Conclusion
This paper demonstrates the feasibility of extending the optimization of Stefan problems to scenarios where the flow in the fluid phase is significant and cannot be ignored. By treating the velocity field as a known variable in the adjoint equations, we were able to derive an incomplete continuous adjoint formulation for the two-phase problem. Our results indicate that the adjoint-based gradient descent method effectively controls the shape of melting fronts in this highly nonlinear system, while greatly reducing the number of function evaluations compared with a derivative-free particle-swarm algorithm. The present formulation is limited to two-dimensional, laminar flows with a prescribed top boundary temperature in a periodic domain and an incomplete adjoint in which the Navier–Stokes adjoint problem is not solved. While the numerical examples indicate that the resulting gradients are accurate enough for practical optimization, the method does not capture sensitivities associated with changes in the flow field itself. This approach nevertheless shows great potential for applications in more complex and dynamic melting processes. Future work could explore more complex actuator basis, non-periodic domains, and three-dimensional systems to further enhance the optimization framework and broaden its applicability.
@@
T.F. and T.S. conceived the study. T.F. performed the numerical simulations. T.F. wrote the original draft with feedback from all authors. T.S. supervised the study.